# Beyond Classical Thermodynamics: Insight into the Microstructures of Black Holes with Quantum-Triggered Violations

### v1

@@ III.2 Thermodynamic curvature via $P$-$V$ criticality
IV $P$-$V$ Criticality and Microstructure of Charged AdS Black Hole with Quantum Anomaly for $\alpha_{\rm c}=0$
@@
IV.1 Coexistence Curves of the $P$-$V$ Phase Transition
@@
Figure 4: $\widetilde{P}$-$\widetilde{V}$ isotherms for the reduced equation of state of charged AdS black hole with quantum anomaly Hu:2024ldp . It shows the oscillatory behaviors for both of $T<T_{\rm c}$ and $T>T_{\rm c}$, indicating the existence of small and large black hole phases which undergo a first-order phase transitions occur both above and below the critical temperature.
@@
$\widetilde{P}(t,\omega)=1+\frac{8}{5}t-\frac{4}{15}t\omega-\frac{8}{135}\omega^{3}+\mathscr{O}(t\omega^{3},\omega^{4})\,.$ (32)
Employing Maxwell’s equal area law as the same method discussed above section, we obtain the non-trivial reduced volumes of the coexistence small and large phases of black hole with quantum anomaly near the critical point, which are
@@
We observe that the small and large black hole phases are positioned on the left and right, respectively, with the dashed curves on either side depicting them and converging at the critical point (marked as a black dot). This coexistence phase is evident for both $T<T_{\rm c}$ and $T>T_{\rm c}$, indicating the presence of small and large black hole phases that undergo first-order phase transitions both above and below the critical temperature.
IV.2 Thermodynamic curvature via $P$-$V$ criticality
Inserting Eqs.(16), (30) and (31) into (III.2), we derive the normalized scalar curvature for the charged AdS black hole incorporating quantum anomaly with $\alpha_{c}={Q^{2}}/{8}$, as obtained below
@@ IV.2 Thermodynamic curvature via $P$-$V$ criticality
Note that $R_{\rm N}$ does not explicitly depend on the black hole charge $Q$–all black holes with different charge share the same expression in the reduced parameter space, a universal result.
In the vicinity of the curve defined by
@@
$R_{\rm N}$ undergoes dramatic changes, approaching negative infinity on this curve. This abrupt change suggests that the microstructure of the black hole undergoes rapid alterations near the temperature $\tilde{T}_{\rm div}$. Furthermore, we calculate the curves where $R_{\rm N}$ changes its sign, given by
@@
indicating a transition between attractive and repulsive interactions within the microstructure as one traverses these curves. These curves are universal, encompassing all charged AdS black holes exhibiting quantum anomalies, irrespective of the specific value of $Q$.
@@
$R_{\rm N}^{s}$ $=$ $\frac{-2-\sqrt{3}}{t^{2}}+\frac{34+20\sqrt{3}}{t}+...O(t^{1},...)\,,$ (38)
$R_{\rm N}^{l}$ $=$ $\frac{-2+\sqrt{3}}{t^{2}}+\frac{34-20\sqrt{3}}{t}+...O(t^{1},...)\,,$ (39)
@@
This result differs from the corresponding value obtained for a van der Waals fluid, which is1/818-1/8- 1 / 8 Wei:2019uqg .
@@
Figure 6: The normalized thermodynamic curvature $R_{\rm N}$ of the charged AdS black hole with quantum anomaly along the coexistence small and large phases. For $T<T_{\rm c}$, the $R_{\rm N}$ changes the sign to positive for the both of small and large black hole phase cases. For $T>T_{\rm c}$, the $R_{\rm N}$ always negative along the coexistence large black hole phase, while the $R_{\rm N}$ can be positive as the temperature divided from critical temperature.
@@
Figure 6 demonstrates that at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity ($R_{\rm N}\rightarrow-\infty$) in both the small and large phases. As the temperature deviates from the critical temperature, $\left|R_{\rm N}\right|$ decreases. It is noteworthy that for temperatures below the critical temperature ($T<T_{\rm c}$), $R_{\rm N}$ changes sign to positive in both the small and large black hole phases, suggesting that a weak repulsive interaction dominates in charged AdS black holes with quantum anomalies at low temperatures (despite similarities to a van der Waals (VdW) fluid in other respects). For temperatures above the critical temperature ($T>T_{\rm c}$), $R_{\rm N}$ remains negative along the coexistence curve for the large black hole phase, indicating that an attractive interaction dominates in charged AdS black holes with quantum anomalies. However, along the coexistence curve for the small black hole phase, $R_{\rm N}$ can be positive as the temperature deviates from the critical temperature, implying that the microstructural interaction in the coexistence small phase of charged AdS black holes with quantum anomalies undergoes a transition from attractive to repulsive during the phase transition.
V Conclusions and Discussion
@@
In this paper, we have utilized the Ruppeiner thermodynamic geometry as a powerful framework to investigate the microscopic structure of charged AdS black holes exhibiting quantum anomalies. Considering quantum effects, we analyzed two scenarios: $\alpha_{\rm c}\leqslant 0$ and $\alpha_{\rm c}=Q^{2}/8$. Furthermore, for both cases, we investigated the volumes corresponding to the coexistence of small and large black holes during the $P$-$V$ phase transitions, and examined the divergence behavior of the thermodynamic curvature scalar $R_{\rm N}$.
@@
For the first scenario ($\alpha_{\rm c}\leqslant 0$), the critical behavior of $R_{\rm N}$ revealed a universal critical exponent of $2$ along both the saturated small and large black hole curves, accompanied by a universal coefficient of $-1/8$. This coefficient aligns precisely with the numerical results obtained for a van der Waals fluid. Notably, at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity in both the small and large phases. As the temperature deviates from the critical temperature, $\left|R_{\rm N}\right|$ decreases, becoming negligible at lower temperatures. In contrast to a van der Waals fluid system, we observed that weak repulsive interactions prevail in charged AdS black holes with quantum anomalies at lower temperatures.
@@
For the second scenario ($\alpha_{\rm c}=Q^{2}/8$), although $R_{\rm N}$ exhibits a universal critical exponent of $2$ along both the small and large black hole curves, its coefficient of $-2\pm\sqrt{3}$ differs significantly from the corresponding value of $-1/8$ obtained for a van der Waals fluid. Moreover, at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity ($R_{\rm N}\rightarrow-\infty$) in both the small and large phases. As the temperature deviates from the critical temperature, $\left|R_{\rm N}\right|$ decreases. It is noteworthy that for temperatures below the critical temperature ($T<T_{\rm c}$), $R_{\rm N}$ changes sign to positive in both the small and large black hole phases, suggesting that a weak repulsive interaction dominates in charged AdS black holes with quantum anomalies at low temperatures (despite similarities to a van der Waals (VdW) fluid in other respects). For temperatures above the critical temperature ($T>T_{\rm c}$), $R_{\rm N}$ remains negative along the coexistence curve for the large black hole phase, indicating that an attractive interaction dominates in charged AdS black holes with quantum anomalies. However, along the coexistence curve for the small black hole phase, $R_{\rm N}$ can be positive as the temperature deviates from the critical temperature, implying that the microstructural interaction in the coexistence small phase of charged AdS black holes with quantum anomalies undergoes a transition from attractive to repulsive during the phase transition.

### v2

@@ III.2 Thermodynamic curvature via $P$-$V$ criticality
IV $P$-$V$ Criticality and Microstructure of Charged AdS Black Hole with Quantum Anomaly for $0<\alpha_{c}<\frac{Q^{2}}{8}$
We restricted our analysis in the previous section to the case where $\alpha_{c}\leqslant 0$. In contrast, accounting for the non-negativity of the central charge and the existence of a critical point narrows the physically viable parameter region to $0<\alpha_{c}\leqslant\frac{Q^{2}}{8}$. Therefore, we focus our subsequent investigations on the phase transitions and microstructure of the black hole in this regime.
IV.1 Coexistence curves of the $P$-$V$ phase transition
In this physical case, in addition to the critical point obtained in (III.1), there is another critical point Hu:2024ldp
$V_{\rm c}$ $=$ $\frac{4\pi}{3}(-12\alpha_{\rm c}+3Q^{2}-K)^{3/2}\,$,
$T_{\rm c}$ $=$ $\frac{3Q^{2}+K}{48\pi\alpha_{\rm c}\sqrt{-12\alpha_{\rm c}+3Q^{2}-K}}\,$,
$P_{\rm c}$ $=$ $\frac{-18\alpha_{\rm c}+6Q^{2}-K}{24\pi(-12\alpha_{\rm c}+3Q^{2}-K)^{2}}\,.$ (31)
It is worth noting that when $\alpha_{\rm c}$ tends to zero, the critical point (IV.1) shows a divergent behavior and is therefore not considered herein, while the critical values presented in Eq. (III.1) reduce to those of the Reissner-Nordström-Anti-de Sitter (RN-AdS) black hole Kubiznak:2012wp . However, when the central charge $\alpha_{\rm c}$ and the electric charge parameter $Q$ satisfy the specific relation $\alpha_{\rm c}=Q^{2}/8$, the auxiliary quantity $K$ (appearing in both Eqs. (III.1) and (IV.1)) becomes zero, and the two critical points merge into one–this intriguing specific case will be investigated in detail in the next section. Therefore, we focus our discussion here on the critical behavior of quantum-anomalous charged AdS black holes with a central charge in the range $0<\alpha_{\rm c}<Q^{2}/8$, which is demonstrated based on Eqs.(III.1), (III.1) and (IV.1) and shown in Fig. 4.
Figure 4: $\widetilde{P}-$$\widetilde{V}$ isotherms for the reduced equation of state of a quantum-anomalous charged AdS black hole. Curves correspond to normalized temperatures $\widetilde{T}=T/T_{\rm c}$: dashed red ($\widetilde{T}=0.8$), dashed blue ($\widetilde{T}=0.9$), dashed black ($\widetilde{T}=1$), solid black ($\widetilde{T}=1$), solid green ($\widetilde{T}=1.1$), solid red ($\widetilde{T}=1.2$). Dashed curves are plotted via Eq.(III.1), while solid curves are via Eq.(IV.1). Parameters are set as $\alpha_{\rm c}=0.1$ and $Q=1$.
In Fig.4, both the dashed and solid curves are derived from the equation of state (III.1): the former uses the critical point in Eq.(III.1), and the latter uses the alternative critical point in Eq.(IV.1). Notably, phase transitions associated with the dashed curves occur only for temperatures below the critical temperature, whereas those corresponding to the solid curves emerge for temperatures above the critical temperature. [^2] This behavior differs from that of van der Waals systems and most black hole systems, where coexisting phases typically appear below the critical temperature Kubiznak:2012wp ; Altamirano:2014tva ; Kubiznak:2016qmn . Additionally, the system exhibits a “thermodynamic singularity” [^3] This singularity manifests as a common intersection point for different isothermal lines in the $\widetilde{P}-$$\widetilde{V}$ diagram, a feature that has also been identified in some previous studies on black hole thermodynamics Frassino:2014pha ; Hennigar:2016ekz ., defined by the temperature independence of pressure, $i.e.$, $(\partial\widetilde{P}/\partial\widetilde{T})_{\widetilde{V}}=0$, see more detailed discussion in Appendix A.
By expanding the equation of state (13) around the critical point (IV.1) via a Taylor series, we obtain the following result
$\widetilde{P}(t,\omega)=1+a_{10}t+a_{11}t\omega+a_{03}\omega^{3}+...\mathscr{O}(t\omega^{3},\omega^{4})$ (32)
where the coefficients $a_{10}$, $a_{11}$, and $a_{03}$ in Eq.(32) are given by Hu:2024ldp as follows
$a_{10}$ $=$ $\frac{4(3Q^{2}-4\alpha_{\rm c}-K)}{9Q^{2}-22\alpha_{\rm c}}\,$, (33)
$a_{11}$ $=$ $\frac{2(5K+3Q^{2}-24\alpha_{\rm c})}{27Q^{2}-66\alpha_{\rm c}}\,,\,\,a_{03}=\frac{40\alpha_{\rm c}-K-15Q^{2}}{27(9Q^{2}-22\alpha_{\rm c})}\,.\,\,$
The variation of these coefficients with $\alpha_{\rm c}$ is visualized in Fig.5.
Figure 5: The coefficients $a_{ij}$ as a function of $\alpha_{\rm c}$ in the range $(0,{Q^{2}}/{8}]$. Here, we set $Q=1$.
Following the same approach as in the previous section, we employ Maxwell’s equal-area law and, from Eq.(32), derive the non-trivial reduced volumes corresponding to the coexisting small and large black hole phases of the quantum-anomalous charged AdS black hole near the critical point defined in Eq.(IV.1). These volumes are given by [^4] The difference between the reduced volumes of the large and small black hole phases $\widetilde{V}_{l}-\widetilde{V}_{s}\propto|t|^{{1}/{2}}$, which gives one of the critical exponents ($1/2$) for the coexistence curve in the context of scaling laws.
$\widetilde{V}_{s}$ $=$ $1-\sqrt{\frac{18K}{3Q^{2}-8\alpha_{\rm c}}(\widetilde{T}-1)}\,$,\,\,\,\,\, (34)
$\widetilde{V}_{l}$ $=$ $1+\sqrt{\frac{18K}{3Q^{2}-8\alpha_{\rm c}}(\widetilde{T}-1)}\,.\,\,$ (35)
To gain a deeper understanding of the phase transition behavior of these coexisting small and large phases in the vicinity of the critical point, we plot the temperature-volume diagram ($\widetilde{T}-$$\widetilde{V}$) based on Eqs. (21), (22), (34) and (35); this diagram is presented in Fig. 6.
Figure 6: Coexistence curves: $\widetilde{T}-$$\widetilde{V}$ phase structure of the quantum-anomalous charged AdS black hole. Here, we set $\alpha_{\rm c}=0.1$ and $Q=1$.
As observed in Fig. 6, the small and large black hole phases are located on the left and right sides of the $\widetilde{T}-$$\widetilde{V}$ diagram, respectively. The blue and red curves on either side correspond to these two phases, with both curves converging at the critical point (denoted by a black dot). Notably, the coexistence of these two phases is evident for both $T<T_{\rm c}$ and $T>T_{\rm c}$, which indicates that the small and large black hole phases undergo first-order phase transitions both below and above the critical temperature.
IV.2 Thermodynamic curvature via $P$-$V$ criticality
By substituting Eqs. (13), (16), and (IV.1) into Eq. (III.2), we derive the normalized scalar curvature $R_{\rm N}$ for the quantum-anomalous charged AdS black hole within the parameter range $0<\alpha_{\rm c}<{Q^{2}}/{8}$. One of the resulting expressions is given as
$R_{\rm N}=\frac{24\alpha_{\rm c}(4\alpha_{\rm c}-2Q^{2}+\eta{\widetilde{V}}^{2/3})[\Theta+12\alpha_{\rm c}(4\alpha_{\rm c}+\eta{\widetilde{V}}^{2/3})]}{[\Theta+24\alpha_{\rm c}(4\alpha_{\rm c}-Q^{2}+\eta{\widetilde{V}}^{2/3})]^{2}}\,\,\,\,\,$ (36)
where the auxiliary quantities $\Theta$ and $\eta$ are defined as
$\Theta$ $=$ $K\widetilde{T}(12\alpha_{\rm c}{\widetilde{V}}^{1/3}-\eta\widetilde{V})-3Q^{2}[(8-12\widetilde{T}{\widetilde{V}}^{1/3})\alpha_{\rm c}+\eta\widetilde{T}\widetilde{V}$]\,,
$\eta$ $=$ $3Q^{2}-12\alpha_{\rm c}-K\,.$
Notably, $R_{\rm N}$ exhibits an explicit dependence on the black hole’s electric charge $Q$ and central charge $\alpha_{\rm c}$. In the low-temperature limit ($\widetilde{T}\rightarrow 0$), this quantity also converges to a constant value of $1/2$. In the vicinity of the curve defined by
$\widetilde{T}_{\rm div}=\frac{24\alpha_{\rm c}[Q^{2}(2-3\widetilde{V}^{2/3})-4\alpha_{\rm c}+\widetilde{V}^{2/3}(K+12\alpha_{\rm c})]}{\widetilde{V}(K^{2}-9Q^{4})+12\alpha_{\rm c}(\widetilde{V}^{1/3}+\widetilde{V})(K+3Q^{2})}\,\,\,\,$ (37)
$R_{\rm N}$ undergoes a dramatic variation, tending to negative infinity exactly on this curve. This abrupt change implies that the microstructure of the black hole undergoes significant alterations near the temperature $\widetilde{T}_{\rm div}$. Furthermore, we calculate the curves where $R_{\rm N}$ changes sign, which are given by
$\widetilde{T}_{0}=\frac{\widetilde{T}_{\rm div}}{2}\,$, (38)
indicating that traversing these curves corresponds to a transition between attractive and repulsive interactions among the black hole’s microstates.
Finally, we discuss the behavior of the thermodynamic curvature as it approaches the critical point along the coexistence curve. To this end, we express the normalized scalar curvature $R_{\rm N}$ as a function of $\widetilde{T}$ by substituting Eqs. (34), (35), and (36); we then derive the series expansions of $R_{\rm N}$ at $\widetilde{T}$ along the coexistence curve for the small and large black hole phases, respectively, which are given by
$R_{\rm N}^{s}$ $=$ $-\frac{1}{8t^{2}}-\frac{[3(Q^{2}-8\alpha_{\rm c})]\xi}{{[(3Q^{2}-8\alpha_{\rm c})]}^{5/4}t^{3/2}}+...\mathscr{O}(t^{-1})\,$,\,\,\, (39)
$R_{\rm N}^{l}$ $=$ $-\frac{1}{8t^{2}}+\frac{[3(Q^{2}-8\alpha_{\rm c})]\xi}{{[(3Q^{2}-8\alpha_{\rm c})]}^{5/4}t^{3/2}}+...\mathscr{O}(t^{-1})\,$,\,\,\, (40)
where $\xi$ is defined as
$\xi=\frac{768\alpha_{\rm c}^{2}+80\alpha_{\rm c}K-31KQ^{2}+63Q^{4}-456\alpha_{\rm c}Q^{2}}{4\sqrt{2}(8\alpha_{\rm c}+K-Q^{2})}\,.$
From these expansions, we observe that $R_{\rm N}$ tends to -$\infty$ at the critical point, with a universal critical exponent of $2$. Furthermore, by substituting Eqs. (39) and (40) and neglecting higher-order terms, we obtain the following limiting expression
$\lim_{t\rightarrow 0}R_{\rm N}t^{2}=-\frac{1}{8}\,.$ (41)
This analysis reveals a universal dimensionless constant of 1/8-1/8, which is in perfect agreement with the numerical results for the van der Waals fluid reported in Wei:2019uqg .
To visualize the divergent behavior of the normalized thermodynamic curvature $R_{\rm N}$ for quantum-anomalous charged AdS black holes (within the parameter range $0<\alpha_{c}<{Q^{2}}/{8}$), we plot $R_{\rm N}$ along the respective coexistence curves of the small and large black hole phases in Fig.7, based on Eqs.(24) and (36).
Figure 7: Normalized thermodynamic curvature $R_{\rm N}$ as a function of $\widetilde{T}=T/T_{\rm c}$ for quantum-anomalous charged AdS black holes along their coexistence phases. The blue and red curves correspond to the small and large black hole phases, respectively; both phases exist for $T<T_{\rm c}$ and $T>T_{\rm c}$, with $R_{\rm N}$ diverging to -$\infty$ as $\widetilde{T}\rightarrow 1$ ($i.e.,T\rightarrow T_{\rm c}$). Here we set $\alpha_{\rm c}=0.1,Q=1$.
Figure 7 demonstrates that at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity ($R_{\rm N}\rightarrow$-$\infty$) in both the small and large phases. It is noteworthy that for temperatures below ($T<T_{\rm c}$) and above ($T>T_{\rm c}$) the critical temperature, $R_{\rm N}$ can be positive as the temperature deviates from the critical temperature, implying that the microstructural interaction in the coexistence small and large phases of charged AdS black holes with quantum anomalies undergoes a transition from attractive to repulsive during the phase transition.
V $P$-$V$ Criticality and Microstructure of Charged AdS Black Hole with Quantum Anomaly for $\alpha_{\rm c}={Q^{2}}/{8}$
@@ IV.2 Thermodynamic curvature via $P$-$V$ criticality
V.1 Coexistence curves of the $P$-$V$ phase transition
@@
Figure 8: $\widetilde{P}-$$\widetilde{V}$ isotherms for the reduced equation of state of charged AdS black hole with quantum anomaly Hu:2024ldp . It shows the oscillatory behaviors for both of $T<T_{\rm c}$ and $T>T_{\rm c}$, indicating the existence of small and large black hole phases which undergo a first-order phase transitions occur both above and below the critical temperature. Here we set $Q=1$.
@@
$\widetilde{P}(t,\omega)=1+\frac{8}{5}t-\frac{4}{15}t\omega-\frac{8}{135}\omega^{3}+...\mathscr{O}(t\omega^{3},\omega^{4})\,.$ (44)
Employing Maxwell’s equal area law as the same method discussed above sections, we obtain the non-trivial reduced volumes of the coexistence small and large phases of black hole with quantum anomaly near the critical point, which are
@@
We observe that the small and large black hole phases are positioned on the left and right, respectively, with the blue and red curves on either side depicting them and converging at the critical point (marked as a black dot). This coexistence phase is evident for both $T<T_{\rm c}$ and $T>T_{\rm c}$, indicating the presence of small and large black hole phases that undergo first-order phase transitions both above and below the critical temperature.
V.2 Thermodynamic curvature via $P$-$V$ criticality
Inserting Eqs.(16), (42) and (43) into (III.2), we derive the normalized scalar curvature for the charged AdS black hole incorporating quantum anomaly with $\alpha_{\rm c}={Q^{2}}/{8}$, as obtained below
@@ V.2 Thermodynamic curvature via $P$-$V$ criticality
Note that $R_{\rm N}$ does not explicitly depend on the black hole charge QQ–all black holes with different charge share the same expression in the reduced parameter space, a universal result. Moreover, as $\widetilde{T}\rightarrow 0$, our analysis reveals that $R_{\rm N}$ approaches the asymptotic value $1/2$. In the proximity of the curve formulated by
@@
$R_{\rm N}$ undergoes dramatic changes, approaching negative infinity on this curve. This abrupt change suggests that the microstructure of the black hole undergoes rapid alterations near the temperature $\widetilde{T}_{\rm div}$. Furthermore, we calculate the curves where $R_{\rm N}$ changes its sign, given by
@@
indicating a transition between attractive and repulsive interactions within the microstructure as one traverses these curves. Notably, these curves exhibit universality, as they apply to all charged AdS black holes (with quantum anomalies) regardless of the specific value of the electric charge $Q$.
@@
[…]
@@
This result differs from the corresponding value obtained for a van der Waals fluid, which is1/8-1/8 Wei:2019uqg ; Wei:2019yvs .
@@
Figure 10: Normalized thermodynamic curvature $R_{\rm N}$ for charged AdS black holes with quantum anomaly along their coexistence phases. The blue curve represents the small black hole phase, and the dashed red curve represents the large black hole phase; both phases exist for $\widetilde{T}=T/T_{\rm c}<1(i.e.,T<T_{\rm c})$ and $\widetilde{T}=T/T_{\rm c}>1(i.e.,T>T_{\rm c})$.
@@
Figure 10 demonstrates that at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity ($R_{\rm N}\rightarrow$-$\infty$) in both the small and large phases. It is noteworthy that for temperatures below the critical temperature ($T<T_{\rm c}$), $R_{\rm N}$ changes sign to positive in both the small and large black hole phases, suggesting that a weak repulsive interaction dominates in charged AdS black holes with quantum anomalies at low temperatures. For temperatures above the critical temperature ($T>T_{\rm c}$), $R_{\rm N}$ remains negative along the coexistence curve for the large black hole phase, indicating that an attractive interaction dominates in charged AdS black holes with quantum anomalies. However, along the coexistence curve for the small black hole phase, $R_{\rm N}$ can be positive as the temperature deviates from the critical temperature, implying that the microstructural interaction in the coexistence small phase of charged AdS black holes with quantum anomalies undergoes a transition from attractive to repulsive during the phase transition.
VI Conclusions and Discussion
@@
In this paper, we adopt Ruppeiner thermodynamic geometry as a robust framework to investigate the microscopic structure of charged AdS black holes with quantum anomalies. Incorporating quantum effects into our analysis, we consider three different parameter scenarios: $\alpha_{\rm c}\leqslant 0$, $0<\alpha_{\rm c}<Q^{2}/8$, and $\alpha_{\rm c}=Q^{2}/8$. For all three scenarios, we further explore the coexistence volumes of small and large black hole phases during $P$-$V$ phase transitions, and systematically examine the divergent behavior of the normalized thermodynamic curvature scalar $R_{\rm N}$.
@@
For the first scenario ($\alpha_{\rm c}\leqslant 0$), the critical behavior of $R_{\rm N}$ revealed a universal critical exponent of $2$ along both small and large black hole curves, accompanied by a universal coefficient of 1/8-1/8. This coefficient aligns precisely with the numerical results obtained for a van der Waals fluid. In particular, at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity in both the small and the large phases. In the limit of $\widetilde{T}\rightarrow 0$, both phases exhibit $R_{\rm N}$ converging to $1/2$, a behavior that reflects their shared asymptotic property at low temperatures. In contrast to a van der Waals fluid system, we observed that weak repulsive interactions prevail in charged AdS black holes with quantum anomalies at lower temperatures.
For the second scenario ($0<\alpha_{\rm c}<Q^{2}/8$), we have further extended our analysis of the phase transition and critical behaviors of the quantum-anomalous charged AdS black hole, and additionally derived the coexistence volumes corresponding to the small and large black hole phases. Furthermore, our investigation of the critical behavior of the normalized scalar curvature $R_{\rm N}$ reveals two key universal features along both the coexistence curves of the small and large black hole phases: a universal critical exponent of $2$, and a universal coefficient of 1/8-1/8. This coefficient is in precise agreement with the numerical results reported for the van der Waals fluid. In particular, $R_{\rm N}$ exhibits divergent behavior in both phases, tending to as the system approaches the critical point. It is also worth emphasizing that for temperatures both below ($T<T_{\rm c}$) and above ($T>T_{\rm c}$) the critical temperature, $R_{\rm N}$ can take positive values when the temperature deviates from the critical point. This observation implies that, during the phase transition of quantum-anomalous charged AdS black holes, the microstructural interactions within the coexisting small and large black hole phases undergo a transition from attractive to repulsive.
@@
For the third scenario ($\alpha_{\rm c}=Q^{2}/8$), although $R_{\rm N}$ exhibits a universal critical exponent of $2$ along both the small and large black hole curves, its coefficient of $2\pm\sqrt{3}$ differs significantly from the corresponding value of 1/8-1/8 obtained for a van der Waals fluid. Moreover, at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity ($R_{\rm N}\rightarrow$-$\infty$) in both the small and large phases. It is noteworthy that for temperatures below the critical temperature ($T<T_{\rm c}$), $R_{\rm N}$ changes sign to positive in both the small and large black hole phases, suggesting that a weak repulsive interaction dominates in charged AdS black holes with quantum anomalies at low temperatures. For temperatures above the critical temperature ($T>T_{\rm c}$), $R_{\rm N}$ remains negative along the coexistence curve for the large black hole phase, indicating that an attractive interaction dominates in charged AdS black holes with quantum anomalies. However, along the coexistence curve for the small black hole phase, $R_{\rm N}$ can be positive as the temperature deviates from the critical temperature, implying that the microstructural interaction in the coexistence small phase of charged AdS black holes with quantum anomalies undergoes a transition from attractive to repulsive during the phase transition.